Divide 56 into two parts such that three times the first part exceeds one third of the second by 48. The parts are ?
Answer by Boreal(15235)

the first part is x
the second part is 56-x.
3x-48 (three times the first exceeds, so subtract 48 to make it equal)=(1/3)(56-x)
3x-48=(1/3)(56-x). Multiply both sides by 3 to remove the fraction
9x-144=56-x
10x=200
x=20, first part.
3 times that is 60
subtract 48 and it is 12
that is 1/3 the second, so 3 times that is the second, and that is 36, second part.
20 and 36 equal 56.